# Designing an Optimal Sensor Network via Minimizing Information Loss

### v1

@@ 5.1 Classical Approaches
Using simulation to determine Bayesian optimal designs dates back to the seminal work of MULLER2005509. The expected utility is approximated pointwise by averaging over samples drawn from the joint distribution of model parameters and observations, given the design. Then, a stochastic optimization algorithm or Monte Carlo sampling can be employed to approximate the solution of the optimization problem (ryan2016review; HUAN2013288). For some utility functions, such as EIG, we require nested Monte Carlo estimators for the pointwise evaluation of the criterion (rainforth2018nesting), which has motivated the use of amortized approaches (rainforth2024modern). In the GP context, popular utilities yield closed-form expressions such as MES or the discrete version of the IMSE (gramacy2020surrogates, Ch. 6). However, optimizing these criteria can still be very challenging, it often requires resorting to sequential and greedy approaches. Related work in mobile sensor networks has employed sequential/adaptive sampling strategies based on classical criteria such as IMSE or MES (xu2011sequential; nguyen2021mobile). Furthermore, another potential difficulty is the need to select designs from a discrete set of possibilities. krause2008near proposes a greedy algorithm for maximizing the EIG over a discrete domain. Recent works from the machine learning community focus on finding suitable acquisition functions (active learning) for GPs (riis2022bayesian). In summary, in BOED we require a model, previously fitted using some initial data, a computable utility/criterion, and a search algorithm. In classical BOED, these elements are usually realized independently, resulting in inefficient algorithms.
@@ 5.2 Variational Approaches
As discussed in Section 2.1, the most popular criterion for BOED is the EIG. However, classical estimators of the EIG require a nested loop for each pointwise evaluation. These computational burdens have motivated a recent trend that aims at “amortizing” the search for the optimal design (rainforth2024modern; foster2019variational). Leveraging variational approximations of the posterior or marginal likelihood, foster2020unified derive variational bounds on the EIG that can be optimized over both variational parameters and designs, hence avoiding the need for nested estimators. More importantly, the computation and optimization of EIG are performed jointly. In this paper, we leverage simulation in a different way, as we do not assume that real data are available to infer a reasonable model. Instead, we may view the MIL criterion as if simulation data arises from a black-box model of the real process, provided by the simulator, for which (hyper)parameters and the ED utility are optimized jointly.
@@ 5.3 Gaussian Process-Based Approaches
Though our application of sparse variational GPs using simulation data to spatiotemporal experimental design is, to our knowledge, novel, it has long been recognized that sparse GP learning appears similar to sensor placement tasks. For example, krause2008near remarks that the original inducing point method of snelson2005sparse is similar and more tractable. More recently, jakkala2023efficient use the sparse variational lower bound, with no data, to geometrically determine inducing points based on the GP kernel.
One related approach using ST-SVGPs, in particular, was proposed by herrera2022spatial, combining the ST-SVGP with the algorithm of krause2008near to create a sensor network with high mutual information in an application to air-quality data. To do so, an ST-SVGP is first trained on a set of “training points;” Krause’s algorithm is subsequently applied to a set of “candidate points,” and the ST-SVGP is retrained. To apply Krause’s algorithm, the mean spatial covariance is used, i.e.,
@@
herrera2022spatial apply their method to a set of air quality sensors in London and find modest increases in the mutual information accordingly.
While herrera2022spatial is successful in creating sensor networks with high mutual information, their setting is fundamentally different from ours in that they assume the existence of real data (i.e., field observations) to train and optimize their networks. Additionally, the application of Krause’s algorithm is not straightforward without identifying a set of discrete “candidate” locations, as computation grows cubically with the number of candidate points. In our experiments, we use baselines modified for continuous optimization accordingly. Furthermore, air temperature presents a different spatiotemporal pattern from air quality: the former shows a robust diurnal cycle, while the latter is largely influenced by point sources.
@@ 6 Case Study: Air Temperature Modeling in Arizona
The motivation for the MIL loss was to effectively utilize large amounts of physics-based simulation data in training a statistical model, which,in turn,is used to identify optimal sensor locations to observe a spatiotemporally varying process. One potential application for our proposed method lies in the climate sciences, where expensive physics-based simulations may be used to create a physically motivated statistical model that enables us to identify optimal sensor locations to observe air temperature, a spatiotemporally varying process. While these physics-based simulations are considered generally accurate, in the sense that they have been compared to historical observations and are physically realistic, they are often not thoroughly evaluated, both in space and time and in terms of the several observables that they produce. As a result, it is valuable to collect observational data from new locations that are maximally informative in order to further evaluate and calibrate the physics-based model. Furthermore, physics-based simulations are often extremely computationally expensive, making it desirable to develop alternative methods to produce climate and weather maps by leveraging sparse but strategically positioned observations. This introduces the experimental design problem of identifying optimal sensor locations. All code used in our experiments is available under an MIT License at https://github.com/DanWaxman/MILSensors.
@@ 6.1 The Physics-Based Model Data
For this application, we rely on simulated data generated by the Weather Research Forecast (WRF) model (powers2017weather). WRF is a physics-based weather and climate model that has been used, among other things, to study heat waves and their interactions with the urban heat island (e.g., li2014quality) as well as the potential of green roofs as heat-mitigating solutions for improving urban resilience (e.g., tewari2019interaction).
@@
Simulations were produced continuously from May 1st (00 UTC) to August 31st (23 UTC) to capture extreme heat conditions and avoid irregular (in space and time) air temperature fluctuations that would result from storm activity typical of the North American monsoon season (vera2006toward). The month of May was considered a spin-up period and was discarded from subsequent analysis. The physics-based model domain was composed of three two-way nested domains centered on Arizona. Here, we focus on the urban grids of the inner domain, specifically the 314 grids encompassing Phoenix, Arizona. This inner domain has a spatial resolution of $2\text{\,}\mathrm{km}$. The multilayer building energy model (MBEM) (salamanca2010new) was used to provide the lower boundary conditions for the coupled WRF atmospheric model. The urban landscape was characterized using three urban categories (i.e., low intensity residential, high intensity residential, and commercial or industrial areas) derived from the US Geological Survey 30-m 2019 National Land Cover Dataset (NLCD) wikham2023thematic. Urban morphology characteristics for each urban category were extracted from burian2002morphological and thermal properties for roofs, roads, and vertical walls were obtained from clarke1991harmonization. A complete description of these urban parameters, including their values, can be found in Table 1 in salamanca2015summertime. Additional information about the physics-based model setup can be found in (salamanca2025effects).
@@ 6.2 MIL-Optimal Points Provide Good Performance
One main hypothesis of our work is that the inducing points in the variational approach will naturally gravitate towards areas that are difficult to predict and, therefore, provide reasonably informative observations, even in the prior kernel. To test this hypothesis, we find the MIL-optimal inducing point locations using data from June 1 to June 30, 2013, and subsequently treat the data from August 1 to August 31, 2013, as ground truth. We then compute the root-mean-squared error (RMSE) in air temperature and the negative predictive log-likelihood (NPLL). For both these measures, lower values indicate better performance; intuitively, RMSE measures how good the mean performances are, and NPLL measures the performance of the predictive distribution.
@@
While we lack “real data” for comparison, the data from June 2013 and August 2013 are quite different, e.g., in terms of maximum daily temperatures, temperature range, etc. We thus argue that utilizing only June 2013 as “simulator data” (i.e., our training data) already imposes a minor effect similar to that of a simulation-to-real data gap.
@@
The other two are model-based approaches popular in spatiotemporal experimental design (mateu2012spatio), based on maximizing the predictive entropy of predictions at the design (jin2012entropy), and minimizing the predictive variance given the design (heuvelink2012sampling); we will refer to these methods as maximum entropy sampling (MES) and integrated minimum square error (IMSE) designs, respectively. Applying these model-based strategies in a fair and consistent manner is not straightforward in our case, since our method has a fundamentally different setting from typical spatiotemporal OED. We discuss the choices we made and the computational improvements in Appendix C.
@@
For all comparisons, we use the kernel hyperparameters learned by the variational GP from the training dataset in the testing dataset. We use a mixture of kernels for the temporal dimension; a Matérn-1/2 kernel for local and non-smooth changes in air temperature, and a Hida-Matérn-3/2 kernel (dowling2021hida) to help capture the diurnal cycle (i.e., daily quasiperiodicity) expected to be present in the data. In the spatial dimension, we use a Matérn-3/2 kernel. We note that while the temporal dimension must be modeled by a Markovian GP, there is no such restriction on the spatial kernel. Indeed, we may even use nonstationary kernels such as the Gibbs kernel (gibbs1998bayesian), which has demonstrated success in modeling climate systems lachland2022kernel. Jointly learning a nonstationary kernel and its inducing point representation introduces its own optimization challenges but does not fundamentally change the theoretical underpinnings of MIL-optimality. We, therefore, include some preliminary experiments using deep kernel learning (wilson2016deep), which learns non-stationary kernels using deep neural networks in Appendix D, but we leave a more detailed study as an interesting direction for future work.
@@
Since in our initial experiment, the largest gains were made with smaller numbers of sensors, we repeated this experiment with 3, 6, 9, 12, and 15 sensors. Results are shown in Fig. 3 and are overall similar to the results of the previous experiments. In particular, the MIL-optimal configuration outperforms random configurations by a considerable margin on average, typically outperforming even the best of 10 uniformly random configurations. In this set of experiments, we additionally compare to MES and IMSE designs. We find that MES performs competitively with a small number of sensors (where space-filling designs are expected to work well); however, MIL becomes the dominant strategy by the time 15 sensors are considered. IMSE performs quite poorly in terms of RMSE, even non-competitively with random and quasi-random design selections — we believe this is part of optimization difficulties, which resulted in designs that do not tend to cover the corners of our domain. In any case, in terms of both RMSE and NPLL, MIL-optimal configurations possess a clear advantage by the time 15 sensors are placed.
@@
Another feature of the MIL-optimal locations is that they have relatively little variance in performance, which suggests some convergence to optimal states. We visualize this in Fig. 5, where the initial states and their corresponding MIL-optimal states are shown. We see that inducing point locations often travel quite far from their initialization points. Interestingly, the random initialization point shown in purple results in a slightly different MIL-optimal location compared to the other two random initializations, but it yields comparable performance. This suggests that there may be multiple slightly different sensor network configurations that are locally MIL-optimal and all perform similarly well. This is apparent in the extremely narrow distribution of RMSE between different initializations in Figs. 3 and 2 when compared to (quasi-)random sensor networks.
@@
On account of the STVGP backbone, MIL-based optimization is extremely efficient. Computing solutions using $N_{t}\cdot N_{s}=720\cdot 314\approx 2\cdot 10^{5}$ data points and $9$ sensor locations takes approximately 5 minutes on modern GPU hardware (Nvidia A100). The computational complexity of the ST-SVGP is $\mathcal{O}(N_{t}M_{s}^{2}+M_{s}^{3})$, thus, for moderate-to-large $T$, we generally expect quadratic scaling with the number of inducing points and linear scaling with respect to the number of spatial and temporal locations (hamelijnck2021spatio). The random and quasi-random designs have trivial computational costs, and MES and IMSE computations take several minutes, even when using Markovian GPs and continuous optimization, as detailed in Appendix C, rather than the typical discrete optimization with kernel-based GPs. We thus conclude that, while MIL is generally more expensive than other OED methods, its cost is on the same order of magnitude as that of other model-based approaches in our case study.
@@ 6.3 Joint Learning is Important
A primary motivation for our method was that greedy learning algorithms, such as BAD may provide strongly suboptimal results in the spatiotemporal setting, while our method instead provides a tractable approach to joint optimization. To visually argue the importance of this joint learning, we optimized the inducing point locations for 9 and 10 sensors separately.
@@
The resulting inducing points, shown in Fig. 4, demonstrate how the optimal locations of inducing points may change when an additional inducing point is added. In other words, the first 9 optimal sensor locations in each network are not the same.
@@ 6.4 Visualizing the Spatiotemporal Performance
In Fig. 8, we visualize the predictions and errors for three randomly selected locations (labeled in Fig. 6 as $1$, 22, and $3$) over three randomly selected days of the test period. We find strong predictive performance and well-calibrated uncertainty estimates. For example, we may consider the miscalibration area, a metric used to assess the calibration of a predictive distribution. For the predicted values shown in Fig. 6. For example, the miscalibration area (chung2021uncertainty) is $0.067$, as shown in Fig. 7. We may also see from Fig. 7 that our predictions tend to be underconfident in this case study, in the sense that our predictive variances are overly conservative.
@@
Also of interest to scientific stakeholders is the number of gross or “extreme” errors in the predictions of a model. For example, maintaining subdegree accuracy, while ultimately an arbitrary line, may be of interest for the scientific validation of a model. We visualize the number of errors greater than 1 degree Celsius as the number of sensors grows from 6 to 15 in Fig. 9, which also visualizes how an MIL-optimal configuration may evolve. It is clear from Fig. 9 that increasing the number of sensor locations greatly reduces the number of extreme errors.

### v2

@@ 5.1 Classical Approaches
Using simulation to determine Bayesian optimal designs dates back to the seminal work of Müller (2005). An expected utility is approximated pointwise by averaging over samples drawn from the joint distribution of model parameters and observations, given the design. Then, a stochastic optimization algorithm or Monte Carlo sampling can be employed to approximate the solution of the optimization problem (Ryan et al., 2016; Huan and Marzouk, 2013). For some utility functions, such as EIG, we require nested Monte Carlo estimators for the pointwise evaluation of the criterion (Rainforth et al., 2018), which has motivated the use of amortized approaches (Rainforth et al., 2024). In the GP context, popular utilities yield closed-form expressions such as MES or the discrete version of the IMSE (Gramacy, 2020, Ch. 6). However, optimizing these criteria can still be very challenging, having to resort to sequential and greedy approaches. Related work in mobile sensor networks has employed sequential/adaptive sampling strategies based on classical criteria such as IMSE or MES (Xu et al., 2011; Nguyen et al., 2021).
Another potential difficulty is the need to select designs from a discrete set of possibilities. Krause et al. (2008) proposes a greedy algorithm for maximizing the EIG over a discrete domain. Recent works from the machine learning community focus on finding suitable acquisition functions (active learning) for GPs (Riis et al., 2022). In summary, in Bayesian OED (BOED) we require a model, previously fitted using some initial data, a computable utility/criterion, and a search algorithm. In classical BOED, these elements are usually realized independently, resulting in inefficient algorithms.
@@ 5.2 Variational Approaches
As discussed in Section 2.1, the most popular criterion for BOED is the EIG. However, classical estimators of the EIG require a nested loop for each pointwise evaluation. These computational burdens have motivated a recent trend that aims at amortizing the search for the optimal design (Rainforth et al., 2024). Leveraging variational approximations of the posterior or the marginal likelihood, Foster et al. (2019, 2020) derive variational bounds on the EIG that can be optimized over both variational parameters and designs, hence avoiding the need for nested estimators. More importantly, the computation and optimization of EIG are performed jointly. In this paper, we leverage simulation in a different way, as we do not assume real data are available to infer a reasonable model. Instead, we may view the MIL criterion as if simulation data arises from a black-box model of the real process, provided by the simulator, for which (hyper)parameters and the utility are optimized jointly.
@@ 5.3 Gaussian Process-Based Approaches
Though our application of sparse variational GPs using simulation data for spatiotemporal experimental design is, to our knowledge, novel, it has long been recognized that sparse GP learning appears similar to sensor placement tasks. For example, Krause et al. (2008) remarks that the original inducing point method of Snelson and Ghahramani (2005) is similar and more tractable. More recently, Jakkala and Akella (2023) use the sparse variational lower bound, with no data, to geometrically determine inducing points based on the GP kernel.
One related approach using ST-SVGPs, in particular, was proposed by Herrera (2022), combining the ST-SVGP with the algorithm of Krause et al. (2008) to create a sensor network with high mutual information in an application to air-quality data. To do so, an ST-SVGP is first trained on a set of “training points;” Krause’s algorithm is subsequently applied to a set of “candidate points,” and the ST-SVGP is retrained. To apply Krause’s algorithm, the mean spatial covariance is used, i.e.,
@@
Herrera (2022) apply their method to a set of air quality sensors in London and find modest increases in the mutual information accordingly.
While Herrera (2022) is successful in creating sensor networks with high mutual information, their setting is fundamentally different from ours in that they assume the existence of real data (i.e., field observations) to train and optimize the networks. Additionally, the application of Krause’s algorithm is not straightforward without identifying a set of discrete “candidate” locations, as computation grows cubically with the number of candidate points. In our experiments, we use baselines modified for continuous optimization accordingly. Furthermore, air temperature presents a different spatiotemporal pattern from air quality: the former shows a robust diurnal cycle, while the latter is largely influenced by point sources.
@@ 6 Case Study: Air Temperature Modeling in Arizona
The motivation for the MIL was to effectively utilize large amounts of physics-based simulation data in training a statistical model, which,in turn,is used to identify optimal sensor locations to observe a spatiotemporally varying process. One potential application for our proposed method lies in the climate sciences, where expensive physics-based simulations may be used to create a physically motivated statistical model that enables us to identify optimal sensor locations to observe air temperature, a spatiotemporally varying process. While these physics-based simulations are considered generally accurate, in the sense that they have been compared to historical observations and are physically realistic, they are often not thoroughly evaluated, both in space and time, and in terms of the several observables that they produce. As a result, it is valuable to collect observational data from new locations that are maximally informative in order to further evaluate and calibrate the physics-based model. Furthermore, physics-based simulations are often extremely computationally expensive, making it desirable to develop alternative methods to produce climate and weather maps by leveraging sparse but strategically positioned observations. This introduces the experimental design problem of identifying optimal sensor locations. All code used in our experiments is available under an MIT License at https://github.com/DanWaxman/MILSensors.
@@ 6.1 The Physics-Based Model Data
For this application, we rely on simulated data generated by the Weather Research Forecast (WRF) model (Powers et al., 2017). WRF is a physics-based weather and climate model that has been used, among other things, to study heat waves and their interactions with the urban heat island (e.g., Li and Bou-Zeid, 2014) as well as the potential of green roofs as heat-mitigating solutions for improving urban resilience (e.g., Tewari et al., 2019).
@@
Simulations were produced continuously from May 1st (00 UTC) to August 31st (23 UTC) to capture extreme heat conditions and avoid irregular (in space and time) air temperature fluctuations that would result from storm activity typical of the North American monsoon season (Vera et al., 2006). The month of May was considered a spin-up period and was discarded from subsequent analysis. The physics-based model domain was composed of three two-way nested domains centered on Arizona. Here, we focus on the urban grids of the inner domain, specifically the 314 grids encompassing Phoenix, Arizona. This inner domain has a spatial resolution of $2\text{\,}\mathrm{km}$. Additional information about the physics-based model setup and its parameterization can be found in Salamanca-Palou et al. (2025) and Salamanca et al. (2015, Table 1), respectively.
@@ 6.2 MIL-Optimal Points Provide Good Performance
One main hypothesis of our work is that the inducing points in the variational approach will naturally gravitate towards areas that are difficult to predict and, therefore, provide reasonably informative observations, even in the prior kernel. To test this hypothesis, we find the MIL-optimal inducing point locations using data from June 1 to June 30, 2013, and subsequently treat the data from August 1 to August 31, 2013, as ground truth. We then compute the root-mean-squared error (RMSE) in air temperature and the negative predictive log-likelihood (NPLL). For both of these measures, lower values indicate better performance; intuitively, RMSE measures how good the mean performance is, and NPLL measures the performance of the predictive distribution.
@@
While we lack “real data” for comparison, the data from June 2013 and August 2013 are quite different, e.g., in terms of maximum daily temperatures, temperature range, etc. We thus argue that utilizing only June 2013 as “simulator data” (i.e., our training data) already imposes a minor effect similar to a simulation-to-real gap.
@@
The other two are model-based approaches popular in spatiotemporal experimental design (Mateu and Müller, 2012), based on maximizing the predictive entropy of predictions at the design (Jin et al., 2012), and minimizing the predictive variance given the design (Heuvelink et al., 2012); we will refer to these methods as maximum entropy sampling (MES) and integrated minimum square error (IMSE) designs, respectively. Applying these model-based strategies in a fair and consistent manner is not straightforward in our case, since our method has a fundamentally different setting from typical spatiotemporal OED. We discuss the choices we made and the computational improvements in Appendix C (Waxman et al., 2025).
@@
For all comparisons, we use the kernel hyperparameters learned by the variational GP from the training dataset in the testing dataset. We use a mixture of kernels for the temporal dimension; a Matérn-1/2 kernel for local and non-smooth changes in air temperature, and a Hida-Matérn-3/2 kernel (Dowling et al., 2021) to help capture the diurnal cycle (i.e., daily quasiperiodicity) expected to be present in the data. In the spatial dimension, we use a Matérn-3/2 kernel. We note that while the temporal dimension must be modeled by a Markovian GP, there is no such restriction on the spatial kernel. Indeed, we may even use nonstationary kernels such as the Gibbs kernel (Gibbs, 1998), which has demonstrated success in modeling climate systems (Lalchand et al., 2022). Jointly learning a nonstationary kernel and its inducing point representation introduces its own optimization challenges but does not fundamentally change the theoretical underpinnings of MIL-optimality. We therefore include some preliminary experiments using deep kernel learning (Wilson et al., 2016), which learns nonstationary kernels using deep neural networks, in Appendix D (Waxman et al., 2025), but we leave a more detailed study as an interesting direction for future work.
@@
Since in our initial experiment, the largest gains were made with smaller numbers of sensors, we repeated this experiment with 3, 6, 9, 12, and 15 sensors. Results are shown in Fig. 3 and are overall similar to the results of the previous experiment. In particular, the MIL-optimal configuration outperforms random configurations by a considerable margin on average, typically outperforming even the best of 10 uniformly random configurations. In this set of experiments, we additionally compare to MES and IMSE designs. We find that MES performs competitively with a small number of sensors (where space-filling designs are expected to work well); however, MIL becomes the dominant strategy by the time 15 sensors are considered. IMSE performs quite poorly in terms of RMSE, even compared to random and quasi-random design selections — we believe this is due to optimization difficulties, which resulted in designs that do not tend to cover the corners of our domain. In any case, in terms of both RMSE and NPLL, MIL-optimal configurations possess a clear advantage by the time 15 sensors are placed.
@@
Another feature of the MIL-optimal locations is their relatively small variance in performance, which suggests some convergence to optimal states. We visualize this in Fig. 5, where the initial states and their corresponding MIL-optimal states are shown. We see that inducing point locations often travel quite far from their initialization points. Interestingly, the random initialization point shown in purple results in a slightly different MIL-optimal location compared to the other two random initializations, but it yields comparable performance. This suggests that there may be multiple slightly different sensor network configurations that are locally MIL-optimal and perform similarly well. This is apparent in the extremely narrow distribution of RMSE between different initializations in Figs. 3 and 2 when compared to (quasi-)random sensor networks.
@@
On account of the STVGP backbone, MIL-based optimization is extremely efficient. Computing solutions using $N_{t}\cdot N_{s}=720\cdot 314\approx 2\cdot 10^{5}$ data points and $9$ sensor locations takes approximately 5 minutes on modern GPU hardware (Nvidia A100). The computational complexity of the ST-SVGP is $\mathcal{O}(N_{t}M_{s}^{2}+M_{s}^{3})$, thus, for moderate-to-large $T$, we generally expect quadratic scaling with the number of inducing points and linear scaling with respect to the number of spatial and temporal locations (Hamelijnck et al., 2021). The random and quasi-random designs have trivial computational costs, and MES and IMSE computations take several minutes, even when using Markovian GPs and continuous optimization, as detailed in Appendix C (Waxman et al., 2025), rather than the typical discrete optimization with kernel-based GPs. We thus conclude that, while MIL is generally more expensive than other OED methods, its cost is on the same order of magnitude as that of other model-based approaches in our case study.
@@ 6.3 Joint Learning is Important
A primary motivation for our method was that greedy learning algorithms, such as BAD may provide strongly suboptimal results in the spatiotemporal setting, while our method instead provides a tractable approach to joint optimization.
@@
To visually argue the importance of this joint learning, we optimized the inducing point locations for 9 and 10 sensors separately. The resulting inducing points, shown in Fig. 4, demonstrate how the optimal locations of inducing points may change when an additional inducing point is added. In other words, the first 9 optimal sensor locations in each network are not the same.
@@ 6.4 Visualizing the Spatiotemporal Performance
In Fig. 8, we visualize the predictions and errors for three randomly selected locations (labeled in Fig. 6 as $1$, 22, and $3$) over three randomly selected days of the test period. We find strong predictive performance and well-calibrated uncertainty estimates. A calibration plot (Fig. 7) shows strong calibration, with a miscalibration area (Chung et al., 2021) of $0.067$. From the calibration plot, we also see that our predictions tend to be underconfident in this case study, in the sense that our predictive variances are overly conservative.
@@
Also of interest to scientific stakeholders is the number of “extreme” errors in the predictions of a model. For example, maintaining subdegree accuracy, while ultimately arbitrary, may be of interest in validating a model. We visualize the number of errors greater than 1 degree Celsius as the number of sensors grows from 6 to 15 in Fig. 9. The results clearly show that increasing the number of sensor locations greatly reduces the number of extreme errors.